OpenStudy (anonymous):

Inverse of the function y = 2x − 6

OpenStudy (anonymous):

@uybuyvf @ganeshie8 @saseal

OpenStudy (anonymous):

make x =???

OpenStudy (anonymous):

or i should say change the x into y then make it back into y = ???

OpenStudy (anonymous):

y=(x+6)/2 Is that it?

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Which of the following expressions represents a function? {(1, 2), (4, –2), (8, 3), (9, –3)} y2 = 16 − x2 2x2 + y2 = 5 x = 7

OpenStudy (anonymous):

@sourwing @Vocaloid

Vocaloid (vocaloid):

the first one, since it assigns a unique output value for each input value

OpenStudy (dinamix):

this invarse of function

OpenStudy (dinamix):

\[f(x)^{-1} =\frac{ x }{ 2 }+3 \]

Vocaloid (vocaloid):

switch f(x) and x, then solve for f(x) x = f(x)/2 + 3 x-3 = f(x)/2 f(x) = 2(x-3)

OpenStudy (abb0t):

@pooja195 can help